OpenStudy (anonymous):

Does -4^-4 simplify to a negative number?

OpenStudy (whpalmer4):

Let's clarify something: is that supposed to be \[(-4)^{-4}\]or\[-(4^{-4})\] ?

OpenStudy (anonymous):

If its the top one then yes. If its the bottom one then no.

OpenStudy (whpalmer4):

@jsz I'll give you a medal if you explain why so @JerryK can learn something from this...

OpenStudy (anonymous):

It's top!

OpenStudy (anonymous):

I'm sorry. The bottom one will simplify to a negative number. -(2)^2=-4 because the negative sign is outside the parentheses. If it were inside the parentheses(-2)^2 =4. It is like multiply by -1 after you square the number. Also remember -2^-4 is like -1/2^4. A negative exponent is a simply a fraction.

OpenStudy (anonymous):

if its the top it will simplify to a positive number.

OpenStudy (whpalmer4):

Yep! If \(a\ne0\), \(a^{-n} = \dfrac{1}{a^n}\) That means our problem can be rewritten as \[(-4)^{-4} = \frac{1}{(-4)^4}\]And because an negative number raised to an even power becomes a positive number, that gives us the answer: a positive number.

OpenStudy (anonymous):

Nice Work. What math are you taking?

OpenStudy (whpalmer4):

I think the last time I took a math class, I was less than half my current age, and it was in the 20th century :-)

OpenStudy (anonymous):

Lol